OpenStudy (anonymous):

write an equation of the line perpendicular to y=1/6x+4that contains(3,-3)?show working

geerky42 (geerky42):

Let say n is slope that is perpendicular to slope m.\[n = -\dfrac{1}{m}\]

geerky42 (geerky42):

So knowing that m is 1/6. Plug in that value then find for n. This is slope for new equation.

geerky42 (geerky42):

Do you need more explanations?

geerky42 (geerky42):

You do know that linear function can be written as y = mx + b, right? Slope of this equation is m. To find another slope equation that is perpendicular to original one. \[n = -\dfrac{1}{m}\] (n is slope for new equation.)